In an hydrogen atom the energy of an electron in the first bohr orbit is -13.12 *105 J/mol.what is the energy required for its excitation to bohr's second orbit?

Grade:12

3 Answers

Nayana Das
7 years ago
Energy of 1st orbit=13.12*10^5(let's say E)
Thus, energy of second orbit=3.28*1-^5(Because E/n^2,let's say E*)
Thus, enery needed to ake from 1st to 2nd orbit=Difference of the two=9.84*10^5 J/mol.
DILIP
7 years ago
The enegry difference between first two orbits of hydrogen atom is the energy required to exite a electron to second orbit from first orbit.
 Energy of 1st orbit = 13.12 x 105 J/mol.
Energy of second orbit = E / n2
Energy of second orbit = 3.28 x 105 j/mol.
Difference btw their energies is 9.84 x 105 J/mol
Study ave
2 years ago
The energy of 1st Bohr’s orbit of H- atom is -13.6 EV. Calculate the energy of the first excited state of Li2+